# SSFlowNet: Semi-supervised Scene Flow Estimation On Point Clouds With Pseudo Label

### v1

@@ 1 Introduction
The emergence of deep neural networks has catalyzed advancements in the field of scene flow estimation. A notable milestone was achieved with the introduction of FlowNet3D Liu et al. (2019), a model that leveraged the principles established by PointNet++ Qi et al. (2017b) to integrate local features into scene flow estimation, thereby pioneering the application of neural networks in this domain. Following this, HPLFlowNet Gu et al. (2019) was developed, featuring an innovative approach that computes multi-scale correlations utilizing upsampling operations within bilateral convolutional layers, thus enhancing the model’s depth perception and contextual understanding. Further advancements were made by Kittenplon et al. (2020), which introduced a novel technique for learning a single step of an unrolled iterative alignededment procedure, thereby refining the accuracy of scene flow predictions. Additionally, the advent of 3DFlow Wang et al. (2022) marked a new era in the field, setting unprecedented benchmarks in both 3D End Point Error (EPE3D) and overall accuracy, thereby exemplifying the dynamic and continuous progression in scene flow estimation methodologies.
@@
Although some work has improved metrics for scene flow estimation, obtaining scene flow labels from the real world can be challenging. Mittal et al. (2019) firstly utilizes an unsupervised method through nearest neighbor loss and cycle consistency loss. However, it ignores the local geometrical properties of point clouds. Inspired by unsupervised methods with point cloud, Wu et al. (2019) further improved the network performance in a coarse-to-fine fashion. However, the errors may accumulate in the early step. Li et al. (2021) proposes a new training mode by pseudo labels, initializing the scene flow through optimal transmission and taking the random walk approach to further improve the label quality. In addition, the pseudo-label-based training approach is also used to address the problem of insufficient label volume. Li et al. (2021) utilized an assignment matrix to guide the generation of pseudo ground truth. Meanwhile, a random walk module was introduced to encourage the local consistency of the pseudo labels. Li et al. (2022a) leveraged monocular images to generate pseudo labels for point clouds, which facilitates the training of scene flow networks. Pseudo-label based training approaches rely on the quality of generated pseudo-labels. Although the pseudo-label based approach has made some progress, the accuracy of pseudo-labels still needs to be improved.
@@
Occupying a unique position in the learning paradigm spectrum, semi-supervised learning (SSL) adeptly mediates the interplay between model accuracy and the magnitude of data. This paper introduces an innovative pseudo-label generation model, a noteworthy advancement in the SSL domain. Grounded in the principles delineated by Puy et al. (2020), our methodology harnesses the capabilities of Graph Neural Networks (GNN) to construct a geometric features graph. This graph is pivotal in developing our correlation matrix, which is meticulously designed to assess the similarities between labeled and unlabeled data points with enhanced precision.
@@ 2 Related Work
Scene Flow Estimation: The concept of scene flow was first articulated by Vedula et al. in Vedula et al. (1999). Scene flow, distinct from the 2D optical flow that delineates the movement trajectories of image pixels, is conceptualized as a vector characterizing the motion of three-dimensional objects. Early research in this field Huguet and Devernay (2007); Menze and Geiger (2015); Vogel et al. (2011) predominantly utilized RGB data. Notably, Huguet and Devernay Huguet and Devernay (2007) introduced a variational approach to estimate scene flow from stereo sequences, while Vogel et al. Vogel et al. (2011) presented a piece-wise rigid scene model for 3D flow estimation. Menze and Geiger Menze and Geiger (2015) advanced the field by proposing an object-level scene flow estimation method, alongside introducing a dataset specifically for 3D scene flow.
The advent of deep learning heralded transformative approaches in scene flow estimation. PointNet Qi et al. (2017a), as a pioneering work, utilized convolutional operations for point cloud feature learning, which was further refined by PointNet++ Qi et al. (2017b) through feature extraction from local domains. Shi and Rajkumar’s Point-GNN Shi and Rajkumar (2020) introduced a graph structure to mitigate translation variance. Building on these foundations, subsequent studies Liu et al. (2019); Puy et al. (2020); Wu et al. (2019); Wang et al. (2022); Li (2023) have achieved impressive results in scene flow estimation. FlowNet3D Liu et al. (2019), for instance, leverages PointNet++ Qi et al. (2017b) for feature extraction and introduces a flow embedding layer to capture and propagate correlations between point clouds for flow estimation. Puy et al. Puy et al. (2020) employed optimal transport for constructing point matches between sequences. Wu et al. Wu et al. (2019) proposed a cost volume module for processing large motions in 3D point clouds, while Wang et al. Wang et al. (2022) innovated an all-to-all flow embedding layer with backward reliability validation to address consistency issues in initial scene flow estimation.
The challenge of acquiring accurate ground truth labels for model training has recently led to the exploration of unsupervised strategies. Mittal et al. Mittal et al. (2019) pioneered this approach in scene flow estimation, employing nearest neighbor loss and cycle consistency loss. Li et al. Li et al. (2021) introduced a pseudo-label generation module that synergizes with the methodology of Mittal et al. Mittal et al. (2019). Furthermore, Li et al. Li et al. (2022b) integrated a local rigidity prior into self-supervised scene flow learning, based on the premise of scenes comprising multiple rigidly moving parts. Ouyang and Raviv Ouyang and Raviv (2021) incorporated an occlusion mask into their data, training the network to detect occlusions more effectively.
@@
Pseudo-Label Training: Contemporary advancements in computer vision heavily rely on deep neural networks. A significant limitation of these networks, however, is their reliance on extensive datasets, typically labeled by human annotators or generated automatically through proxy tasks. Lee et al. Lee et al. (2013) introduced a method for predicting pseudo-labels of unlabeled data by analyzing relationships within dataset samples. In this method, for data without labels, the network selects the class with the maximum predicted probability, a technique analogous to Entropy Regularization. Iscen et al. Iscen et al. (2019) furthered this approach by introducing label propagation, utilizing existing ground truth labels to generate new labels.
In the realm of scene flow estimation, several methods based on pseudo-labels have emerged Shen et al. (2023); Dong et al. (2022); Li et al. (2022a). Shen et al. Shen et al. (2023) employed a self-supervised approach using pseudo-labels, drawing inspiration from Puy et al. Puy et al. (2020) to utilize optimal transport for initializing flow and a random walk technique for label completion. Dong et al. Dong et al. (2022) proposed a method integrating flow into multi rigid-body motion for higher-level scene abstraction. In contrast, Li et al. Li et al. (2022a) developed a multi-modality framework combining RGB images and point clouds to generate pseudo labels for training networks.
@@
Memory in Scene Flow Estimation: Memory modules have become increasingly prevalent in 2D video semantic segmentation frameworks. Traditional semantic segmentation on single frames often overlooks the temporal dimension, leading to inaccuracies, particularly in the context of occlusion. Paul et al. Paul et al. (2021) innovatively proposed the use of local Memory Attention Networks, which effectively transform existing single-frame semantic segmentation models into robust video semantic segmentation frameworks. Similarly, Li et al. Li et al. (2023) introduced a framework for semantic segmentation of temporal sequences in point clouds, utilizing a memory network for the storage and retrieval of past information.
@@ 3.3.1 Flow-Graph Encoder
Most current point cloud feature encoder follow the set abstraction layers of PointNet++ Qi et al. (2017b) called Flow-Graph encoder, which obtains point cloud context by local grouping and max pooling.But it fails to utilize the spatial distribution of the input point cloud. This is because hierarchical feature learning fails to encode the spatial distribution in the division of the point clouds.
@@
In Equation 3, $E$ denotes the edge set in the graph structure. We select the neighboring points around each point and use their distance and feature differences as the criteria for edge formation. ||||| | is the contact operator. Following the methodology suggested by Kittenplon et al. Kittenplon et al. (2020), we encode edge features using three consecutive $setconv$ layers as our convolution mechanism and apply the furthest point sampling method for down-sampling.
In contrast to previous GNN methods, we introduce the concept of spatial memory in the feature extractor. The concept of memory is applied to the field of semantic segmentation Tokmakov et al. (2017); Nilsson and Sminchisescu (2018), study shows that sequential input performs better than one-frame input, because sequential input allows the neural network to learn temporal information. The memory algorithm bases on the assumption of a certain level of temporal consistency from one frame to the next, previous work use GRU(Gate Recurrent Unit) to assist in semantic segmentation or other tasks by carrying prior information through the hidden layer.
@@ 3.3.2 Correlation Matrix
Neighbouring points share similar motion patterns, in the current studies, Kittenplon et al. (2020); Wu et al. (2019) introduces smooth loss for penalising flow vectors that deviate too much from the surrounding points, and Shi and Ma (2022) also combines the knowledge of semantic segmentation with scene flow estimation model with the aim of segmenting objects in the scene that have the same motion model.
@@ 3.4.1 Chamfer Loss
chamfer loss measures the dissimilarity or discrepancy between two sets of points, often used for tasks like point cloud registration, shape matching, and generative modeling. Following Kittenplon et al. (2020); Wu et al. (2019), we calculate wrapped point by $\hat{P}=P+F$, and loss value. Equation shown in equation 7.
@@ 3.4.2 Weighted Smooth Loss
Neighbouring point clouds share flow consistency, and the smooth loss is used to measure the degree of similarity with neighbouring points flow vectors. Smooth loss has been widely used in scene flow estimation Puy et al. (2020); Wu et al. (2019), however, assigning the same weight to all surrounding points may mentor the accumulation of streaming errors.
@@ 4 Experiment
In this section, we train and evaluate our SSLFlowNet on two datasets, FlyingThings3D Mayer et al. (2016) and KITTI Scene Flow Menze et al. (2018). Firstly, we compare with the current mainstream fully and semi-supervised models, as well as the pseudo-label generation model to illustrate the quality of our pseudo-label generation. Then, we combine the pseudo-label generation model with 3DFlownet Wang et al. (2022) and take 1/8,1/16,1/32,1/64 in every batch for auxiliary training. Finally, we perform ablation experiments on different modules, and the experimental results prove that our proposed network structure can optimise the process of flow up-sampling.
@@ 4.1 Experimental Setting
Dataset We conduct our experiments on FlyingThings3D and KITTI respectively. FlyingThings3D Mayer et al. (2016) is a synthetic dataset for optical flow, disparity and scene flow estimation. It consists of everyday objects flying along randomized 3D trajectories. It contains 19640 training examples and 3824 test examples. As most papers do, we only use one-quarter of the training data (4910 pairs). KITTI is a real-world scene flow dataset with 200 pairs for which 142 are used for testing without any fine-tuning. As most scene flow model do, we take the processing in HPLFLownet Gu et al. (2019) to generate the non-occluded datasets. To be fair, we compare with other models on the same dataset as well as in training mode.
@@ 4.2 Evaluation of Pseudo Labels
Our experiment results is shown in table 1. To reflect the generalisation ability of the model, we use random sampling to select 1/8,1/16,1/32,1/64 of the data. Subsequently, we adopt label propagation to spread the ground truth labels to the global. We adopt the same approach as Iscen et al. (2019) and use four metrics to measure the quality of our labels. To be fair, we exclude known labels when calculating EPE and accuracy. Our pseudo-label generation model was trained on Flyingthings for a total of 200 epochs, taking 10 hours. We tested the model on the KITTI dataset without fine-tuning.
Li et al. (2021) uses a consistent approach with Puy et al. (2020) to generate labels, and Iscen et al. (2019) generate pseudo labels by optical flow learning. In contrast with these two, our label generation model is able to achieve higher accuracy. We are also able to generate high-quality pseudo-labels stably under different ground truth data volumes. Our results are shown in Figure 4.
@@ 4.3 Results
We combine the generated pseudo-labels with 3DFlowNet Gu et al. (2019) to show that training by pseudo-labelling is effective. We train on Flythings3D, test model accuracy on the Flythings3D test set and KITTI as shown in Table. 2, and compare it to the original benchmark model. The experimental results show that we sacrifice only a small portion of the accuracy, but greatly reduce the labelling cost.
@@ 4.4 Ablation Study
For our weighted smooth loss experiments, to be fair, we take the averaged smooth loss used in Puy et al. (2020) and Wu et al. (2019). The result was shown in Table. 3.
@@ 5 Discussion
Noise in Point Clouds: Point clouds inherently possess noise or outliers, often attributable to sensor inaccuracies or environmental factors. The noise, an inevitable byproduct of the scanning and reconstruction processes, poses significant challenges in feature extraction and can mislead the search for corresponding points. While several approaches Rakotosaona et al. (2020); Luo and Hu (2020, 2021) have explored denoising via neural networks, effectively mitigating the errors induced by noise points within the context of scene flow estimation remains an open problem.
Predominantly, scene flow models are trained using the simulated dataset FlyingThings3D. However, when these models are transitioned to real-world scenarios such as those represented by KITTI, they exhibit varying degrees of performance degradation. Li et al. Li et al. (2022a) proposed a noisy-label-aware training scheme that leverages the geometric relations of points, yet this approach continues to grapple with limitations in accuracy.
Occlusions and Scene Flow: Occlusions, resulting from the time interval between two frames, can lead to the disappearance of objects and significantly impact flow estimation accuracy. Recent research efforts Saxena et al. (2019); Ouyang and Raviv (2021); Zanfir and Sminchisescu (2015) have focused on developing methodologies to address occlusions. Notably, Ouyang et al. Ouyang and Raviv (2021) introduced a novel framework to assess occlusion in 3D scene flow estimation on point clouds. However, the accuracy of unsupervised and semi-supervised models in predicting occlusions remains suboptimal.
Current models frequently employ the data processing strategy outlined in HPLFlowNet Gu et al. (2019), which simplifies the flow estimation task by eliminating occluded points. This, however, can adversely affect model generalization. Thus, the development of effective pseudo-labels for training models, with the aim of enhancing generalization, presents a promising avenue for future research.

### v2

@@ 1 Introduction
The emergence of deep neural networks has catalyzed advancements in the field of scene flow estimation. A notable milestone was achieved with the introduction of FlowNet3D Liu et al. [2019], a model that leveraged the principles established by PointNet++ Qi et al. [2017b] to integrate local features into scene flow estimation, thereby pioneering the application of neural networks in this domain. Following this, HPLFlowNet Gu et al. [2019] was developed, featuring an innovative approach that computes multi-scale correlations utilizing upsampling operations within bilateral convolutional layers, thus enhancing the model’s depth perception and contextual understanding. Further advancements were made by Kittenplon et al. [2020], which introduced a novel technique for learning a single step of an unrolled iterative alignededment procedure, thereby refining the accuracy of scene flow predictions. Additionally, the advent of 3DFlow Wang et al. [2022] marked a new era in the field, setting unprecedented benchmarks in both 3D End Point Error (EPE3D) and overall accuracy, thereby exemplifying the dynamic and continuous progression in scene flow estimation methodologies.
@@
Although some work has improved metrics for scene flow estimation, obtaining scene flow labels from the real world can be challenging. Mittal et al. [2019] firstly utilizes an unsupervised method through nearest neighbor loss and cycle consistency loss. However, it ignores the local geometrical properties of point clouds. Inspired by unsupervised methods with point cloud, Wu et al. [2019] further improved the network performance in a coarse-to-fine fashion. However, the errors may accumulate in the early step. Li et al. [2021] proposes a new training mode by pseudo labels, initializing the scene flow through optimal transmission and taking the random walk approach to further improve the label quality. In addition, the pseudo-label-based training approach is also used to address the problem of insufficient label volume. Li et al. [2021] utilized an assignment matrix to guide the generation of pseudo ground truth. Meanwhile, a random walk module was introduced to encourage the local consistency of the pseudo labels. Li et al. [2022a] leveraged monocular images to generate pseudo labels for point clouds, which facilitates the training of scene flow networks. Pseudo-label based training approaches rely on the quality of generated pseudo-labels. Although the pseudo-label based approach has made some progress, the accuracy of pseudo-labels still needs to be improved.
@@
Occupying a unique position in the learning paradigm spectrum, semi-supervised learning (SSL) adeptly mediates the interplay between model accuracy and the magnitude of data. This paper introduces an innovative pseudo-label generation model, a noteworthy advancement in the SSL domain. Grounded in the principles delineated by Puy et al. [2020], our methodology harnesses the capabilities of Graph Neural Networks (GNN) to construct a geometric features graph. This graph is pivotal in developing our correlation matrix, which is meticulously designed to assess the similarities between labeled and unlabeled data points with enhanced precision.
@@ 2 Related Work
Scene Flow Estimation: The concept of scene flow was first articulated by Vedula et al. in Vedula et al. [1999]. Scene flow, distinct from the 2D optical flow that delineates the movement trajectories of image pixels, is conceptualized as a vector characterizing the motion of three-dimensional objects. Early research in this field Huguet and Devernay [2007], Menze and Geiger [2015], Vogel et al. [2011] predominantly utilized RGB data. Notably, Huguet and Devernay Huguet and Devernay [2007] introduced a variational approach to estimate scene flow from stereo sequences, while Vogel et al. Vogel et al. [2011] presented a piece-wise rigid scene model for 3D flow estimation. Menze and Geiger Menze and Geiger [2015] advanced the field by proposing an object-level scene flow estimation method, alongside introducing a dataset specifically for 3D scene flow.
The advent of deep learning heralded transformative approaches in scene flow estimation. PointNet Qi et al. [2017a], as a pioneering work, utilized convolutional operations for point cloud feature learning, which was further refined by PointNet++ Qi et al. [2017b] through feature extraction from local domains. Shi and Rajkumar’s Point-GNN Shi and Rajkumar [2020] introduced a graph structure to mitigate translation variance. Building on these foundations, subsequent studies Liu et al. [2019], Puy et al. [2020], Wu et al. [2019], Wang et al. [2022], Li [2023] have achieved impressive results in scene flow estimation. FlowNet3D Liu et al. [2019], for instance, leverages PointNet++ Qi et al. [2017b] for feature extraction and introduces a flow embedding layer to capture and propagate correlations between point clouds for flow estimation. Puy et al. Puy et al. [2020] employed optimal transport for constructing point matches between sequences. Wu et al. Wu et al. [2019] proposed a cost volume module for processing large motions in 3D point clouds, while Wang et al. Wang et al. [2022] innovated an all-to-all flow embedding layer with backward reliability validation to address consistency issues in initial scene flow estimation.
The challenge of acquiring accurate ground truth labels for model training has recently led to the exploration of unsupervised strategies. Mittal et al. Mittal et al. [2019] pioneered this approach in scene flow estimation, employing nearest neighbor loss and cycle consistency loss. Li et al. Li et al. [2021] introduced a pseudo-label generation module that synergizes with the methodology of Mittal et al. Mittal et al. [2019]. Furthermore, Li et al. Li et al. [2022b] integrated a local rigidity prior into self-supervised scene flow learning, based on the premise of scenes comprising multiple rigidly moving parts. Ouyang and Raviv Ouyang and Raviv [2021] incorporated an occlusion mask into their data, training the network to detect occlusions more effectively.
@@
Pseudo-Label Training: Contemporary advancements in computer vision heavily rely on deep neural networks. A significant limitation of these networks, however, is their reliance on extensive datasets, typically labeled by human annotators or generated automatically through proxy tasks. Lee et al. Lee et al. [2013] introduced a method for predicting pseudo-labels of unlabeled data by analyzing relationships within dataset samples. In this method, for data without labels, the network selects the class with the maximum predicted probability, a technique analogous to Entropy Regularization. Iscen et al. Iscen et al. [2019] furthered this approach by introducing label propagation, utilizing existing ground truth labels to generate new labels.
In the realm of scene flow estimation, several methods based on pseudo-labels have emerged Shen et al. [2023], Dong et al. [2022], Li et al. [2022a]. Shen et al. Shen et al. [2023] employed a self-supervised approach using pseudo-labels, drawing inspiration from Puy et al. Puy et al. [2020] to utilize optimal transport for initializing flow and a random walk technique for label completion. Dong et al. Dong et al. [2022] proposed a method integrating flow into multi rigid-body motion for higher-level scene abstraction. In contrast, Li et al. Li et al. [2022a] developed a multi-modality framework combining RGB images and point clouds to generate pseudo labels for training networks.
@@
Memory in Scene Flow Estimation: Memory modules have become increasingly prevalent in 2D video semantic segmentation frameworks. Traditional semantic segmentation on single frames often overlooks the temporal dimension, leading to inaccuracies, particularly in the context of occlusion. Paul et al. Paul et al. [2021] innovatively proposed the use of local Memory Attention Networks, which effectively transform existing single-frame semantic segmentation models into robust video semantic segmentation frameworks. Similarly, Li et al. Li et al. [2023] introduced a framework for semantic segmentation of temporal sequences in point clouds, utilizing a memory network for the storage and retrieval of past information.
@@ 3.3.1 Flow-Graph Encoder
Most current point cloud feature encoder follow the set abstraction layers of PointNet++ Qi et al. [2017b] called Flow-Graph encoder, which obtains point cloud context by local grouping and max pooling.But it fails to utilize the spatial distribution of the input point cloud. This is because hierarchical feature learning fails to encode the spatial distribution in the division of the point clouds.
@@
In Equation 3, $E$ denotes the edge set in the graph structure. We select the neighboring points around each point and use their distance and feature differences as the criteria for edge formation. ||||| | is the contact operator. Following the methodology suggested by Kittenplon et al. Kittenplon et al. [2020], we encode edge features using three consecutive $setconv$ layers as our convolution mechanism and apply the furthest point sampling method for down-sampling.
In contrast to previous GNN methods, we introduce the concept of spatial memory in the feature extractor. The concept of memory is applied to the field of semantic segmentation Tokmakov et al. [2017], Nilsson and Sminchisescu [2018], study shows that sequential input performs better than one-frame input, because sequential input allows the neural network to learn temporal information. The memory algorithm bases on the assumption of a certain level of temporal consistency from one frame to the next, previous work use GRU(Gate Recurrent Unit) to assist in semantic segmentation or other tasks by carrying prior information through the hidden layer.
@@ 3.3.2 Correlation Matrix
Neighbouring points share similar motion patterns, in the current studies, Kittenplon et al. [2020], Wu et al. [2019] introduces smooth loss for penalising flow vectors that deviate too much from the surrounding points, and Shi and Ma [2022] also combines the knowledge of semantic segmentation with scene flow estimation model with the aim of segmenting objects in the scene that have the same motion model.
@@ 3.4.1 Chamfer Loss
chamfer loss measures the dissimilarity or discrepancy between two sets of points, often used for tasks like point cloud registration, shape matching, and generative modeling. Following Kittenplon et al. [2020], Wu et al. [2019], we calculate wrapped point by $\hat{P}=P+F$, and loss value. Equation shown in equation 7.
@@ 3.4.2 Weighted Smooth Loss
Neighbouring point clouds share flow consistency, and the smooth loss is used to measure the degree of similarity with neighbouring points flow vectors. Smooth loss has been widely used in scene flow estimation Puy et al. [2020], Wu et al. [2019], however, assigning the same weight to all surrounding points may mentor the accumulation of streaming errors.
@@ 4 Experiment
In this section, we train and evaluate our SSLFlowNet on two datasets, FlyingThings3D Mayer et al. [2016] and KITTI Scene Flow Menze et al. [2018]. Firstly, we compare with the current mainstream fully and semi-supervised models, as well as the pseudo-label generation model to illustrate the quality of our pseudo-label generation. Then, we combine the pseudo-label generation model with 3DFlownet Wang et al. [2022] and take 1/8,1/16,1/32,1/64 in every batch for auxiliary training. Finally, we perform ablation experiments on different modules, and the experimental results prove that our proposed network structure can optimise the process of flow up-sampling.
@@ 4.1 Experimental Setting
Dataset We conduct our experiments on FlyingThings3D and KITTI respectively. FlyingThings3D Mayer et al. [2016] is a synthetic dataset for optical flow, disparity and scene flow estimation. It consists of everyday objects flying along randomized 3D trajectories. It contains 19640 training examples and 3824 test examples. As most papers do, we only use one-quarter of the training data (4910 pairs). KITTI is a real-world scene flow dataset with 200 pairs for which 142 are used for testing without any fine-tuning. As most scene flow model do, we take the processing in HPLFLownet Gu et al. [2019] to generate the non-occluded datasets. To be fair, we compare with other models on the same dataset as well as in training mode.
@@ 4.2 Evaluation of Pseudo Labels
Our experiment results is shown in table 1. To reflect the generalisation ability of the model, we use random sampling to select 1/8,1/16,1/32,1/64 of the data. Subsequently, we adopt label propagation to spread the ground truth labels to the global. We adopt the same approach as Iscen et al. [2019] and use four metrics to measure the quality of our labels. To be fair, we exclude known labels when calculating EPE and accuracy. Our pseudo-label generation model was trained on Flyingthings for a total of 200 epochs, taking 10 hours. We tested the model on the KITTI dataset without fine-tuning.
Li et al. [2021] uses a consistent approach with Puy et al. [2020] to generate labels, and Iscen et al. [2019] generate pseudo labels by optical flow learning. In contrast with these two, our label generation model is able to achieve higher accuracy. We are also able to generate high-quality pseudo-labels stably under different ground truth data volumes. Our results are shown in Figure 4.
@@ 4.3 Results
We combine the generated pseudo-labels with 3DFlowNet Gu et al. [2019] to show that training by pseudo-labelling is effective. We train on Flythings3D, test model accuracy on the Flythings3D test set and KITTI as shown in Table. 2, and compare it to the original benchmark model. The experimental results show that we sacrifice only a small portion of the accuracy, but greatly reduce the labelling cost.
@@ 4.4 Ablation Study
For our weighted smooth loss experiments, to be fair, we take the averaged smooth loss used in Puy et al. [2020] and Wu et al. [2019]. The result was shown in Table. 3.
@@ 5 Discussion
Noise in Point Clouds: Point clouds inherently possess noise or outliers, often attributable to sensor inaccuracies or environmental factors. The noise, an inevitable byproduct of the scanning and reconstruction processes, poses significant challenges in feature extraction and can mislead the search for corresponding points. While several approaches Rakotosaona et al. [2020], Luo and Hu [2020, 2021] have explored denoising via neural networks, effectively mitigating the errors induced by noise points within the context of scene flow estimation remains an open problem.
Predominantly, scene flow models are trained using the simulated dataset FlyingThings3D. However, when these models are transitioned to real-world scenarios such as those represented by KITTI, they exhibit varying degrees of performance degradation. Li et al. Li et al. [2022a] proposed a noisy-label-aware training scheme that leverages the geometric relations of points, yet this approach continues to grapple with limitations in accuracy.
Occlusions and Scene Flow: Occlusions, resulting from the time interval between two frames, can lead to the disappearance of objects and significantly impact flow estimation accuracy. Recent research efforts Saxena et al. [2019], Ouyang and Raviv [2021], Zanfir and Sminchisescu [2015] have focused on developing methodologies to address occlusions. Notably, Ouyang et al. Ouyang and Raviv [2021] introduced a novel framework to assess occlusion in 3D scene flow estimation on point clouds. However, the accuracy of unsupervised and semi-supervised models in predicting occlusions remains suboptimal.
Current models frequently employ the data processing strategy outlined in HPLFlowNet Gu et al. [2019], which simplifies the flow estimation task by eliminating occluded points. This, however, can adversely affect model generalization. Thus, the development of effective pseudo-labels for training models, with the aim of enhancing generalization, presents a promising avenue for future research.